OpenStudy (anonymous):

What is the molar mass of a gas if 1.30g of the gas has a volume of 245mL at STP?

OpenStudy (anonymous):

start with PV=nRT. Solve for n. n = PV/(RT) stp means 273.15 K and 1 atm. (Must use absolute temperature). R = gas constant. n= (1 atm * .245 )/(273.15 * .08206) = 0.0109303 mols

OpenStudy (anonymous):

molar mass is grams/mole. 1.3g / 0.0109303 mol = 118.9 g/mol